# TomoGraphView: 3D Medical Image Classification with Omnidirectional Slice Representations and Graph Neural Networks

### v1

@@ 6 Conclusion
In this work, we present TomoGraphView, a unified framework for 33D medical image classification that combines omnidirectional volume slicing with spherical graph-based feature aggregation. Using six diverse datasets spanning different anatomical regions, imaging modalities, and classification tasks, we demonstrated that (I) our omnidirectional slicing approach consistently outperforms standard slicing strategies based on axial, sagittal, coronal, or combined views, and (II) spherical graph-based feature aggregation surpasses the performance of existing slice-wise feature aggregation baselines.
Although three-dimensional models hold great promise in medical image analysis, particularly for 3D tomography data, as they leverage spatial relationships across all axes to better capture anatomical context, their adoption in practice is impeded by several challenges: They are computationally demanding, require substantially larger datasets to achieve robust generalization due to their high parameter counts, and offer limited possibilities for transfer learning, since publicly available pretrained 33D networks remain comparatively scarce and less established than 22D alternatives [van2025foundation]. As a consequence, slice-based 22D approaches remain an attractive choice, as they are more efficient to train, naturally benefit from decomposing volumetric scans into multiple slices, and can exploit powerful pretrained backbones developed for natural image tasks.
This is where TomoGraphView comes into the picture. We promote our presented omnidirectional volume slicing strategy as an attractive alternative to existing, canonical volume slicing methods, as it has been shown to improve downstream performance across a plethora of prediction heads, including MLP, LSTM, and transformer-based approaches. Moreover, by integrating omnidirectional volume slicing with spherical graph-based feature aggregation, our TomoGraphView framework achieves superior performance levels when compared against both existing slice-wise feature aggregation methods and large-scale pretrained 33D models. The experimental evaluations with their corresponding results in this work underscore the added value of combining 22D foundation models with advanced volume slicing strategies and feature-aggregation techniques, and therefore represent a key technology toward bridging the gap until fully native 33D foundation models become available for medical image analysis, thereby highlighting the need of our proposed approach to enable more accurate, personalized, and scalable solution for tomography-based image diagnostics.
In summary, our presented work highlights the versatility and promise of 22D slice-wise strategies for 33D medical image classification. We believe that combining 22D foundation models with advanced volumetric slicing techniques and graph-based slice-wise feature aggregation methods represents a key step toward bridging the gap until fully native 33D foundation models become available in medical image analysis.
@@ Acknowledgments
Authors JK, SMF, and JAS gratefully acknowledge financial funding received by Munich Center for Machine Learning, sponsored by the German Federal Ministry of Education and Research. Authors JCP and JK gratefully acknowledge funding from the Wilhelm Sander Foundation in Cancer Research (2022.032.1). Author SMF has received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 515279324 / SPP 2177.
@@
All data, except the in-house soft-tissue sarcoma dataset, used in this study, are publicly available. We provide detailed descriptions and proper citations for each dataset used to ensure full transparency and reproducibility. Additionally, all code for data preprocessing and preparation is publicly released and can be used to fully reproduce our results. The code is available at: http://github.com/compai-lab/2025-MedIA-kiechle. The in-house soft-tissue sarcoma data may be obtained from the corresponding author on reasonable request and may be published online in the future after approval by the institutional ethics review boards. We publicly share our accessible code base at http://github.com/compai-lab/2025-MedIA-kiechle and provide a user-friendly library for omnidirectional volume slicing at https://pypi.org/project/OmniSlicer.
@@ Appendix E Slice-wise Feature Aggregation Methods
Appendix F 33D Backbones Results
@@
TomoGraphView (ours) 0.6999 0.6464 0.6711 0.2924
Kidney CT FMCIB 0.6828 0.6392 0.6785 0.2706

### v2

@@ 6 Conclusion
In this work, we present TomoGraphView, a unified framework for 33D medical image classification that combines omnidirectional volume slicing with spherical graph-based feature aggregation. Using six diverse datasets spanning multiple anatomical regions, imaging modalities, and clinical tasks, we demonstrated that (I) our omnidirectional slicing approach consistently outperforms standard slicing strategies based on axial, sagittal, coronal, or combined views, and (II) spherical graph-based feature aggregation improves performance beyond existing slice-wise feature aggregation baselines.
While 33D models are naturally suited for tomographic data as they capture spatial context across all axes natively, their use in practice remains challenging: they require substantial computational resources, substantially larger datasets to achieve robust generalization due to their high parameter counts, and offer limited possibilities for transfer learning, as publicly available pretrained 33D networks remain comparatively scarce and less established than 22D alternatives [van2025foundation]. As a result, slice-based 22D methods remain an appealing alternative: they are more efficient to train, benefit from the large number of slices extracted from each volume, and can directly leverage powerful 22D foundation models developed for natural images.
TomoGraphView builds on these advantages. Our omnidirectional slicing strategy provides a flexible and effective alternative to standard slicing along canonical planes, improving downstream performance across a wide range of prediction heads, including MLPs, LSTMs, and transformer-based approaches. Moreover, by integrating omnidirectional volume slicing with spherical graph-based feature aggregation, our TomoGraphView framework achieves superior performance levels when compared against both existing slice-wise feature aggregation methods and large-scale pretrained 33D models. The results across our experimental evaluations underscore the value of combining powerful 22D encoders with expressive methods capable of sampling and integrating versatile information from volumetric scans. Therefore, TomoGraphView offers a practical and powerful approach as the field continues to progress toward fully native 3D foundation models.
In summary, this work highlights both the relevance and flexibility of 22D slice-based strategies for 33D medical image classification. By integrating powerful 22D vision foundation models with improved volumetric slicing and graph-based aggregation, TomoGraphView contributes a key step toward bridging the gap until robust and widely available 33D foundation models become standard in medical image analysis.
@@ Acknowledgments
Authors JK, SMF, CIB, and JAS gratefully acknowledge financial funding received by Munich Center for Machine Learning, sponsored by the German Federal Ministry of Education and Research. Authors JCP and JK gratefully acknowledge funding from the Wilhelm Sander Foundation in Cancer Research (2022.032.1). Author SMF has received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 515279324 / SPP 2177.
@@
All data, except the in-house soft-tissue sarcoma dataset, used in this study, are publicly available. We provide detailed descriptions and proper citations for each dataset used to ensure full transparency and reproducibility. Additionally, all code for data preprocessing and preparation is publicly released and can be used to fully reproduce our results. The code is available at http://github.com/compai-lab/2025-MedIA-kiechle. The in-house soft-tissue sarcoma data may be obtained from the corresponding author on reasonable request and may be published online in the future after approval by the institutional ethics review boards. Furthermore, we provide a user-friendly library for omnidirectional volume slicing at https://pypi.org/project/OmniSlicer.
@@ Appendix E Slice-wise Feature Aggregation Methods
Appendix F 33D Pretrained Backbones Results
@@
TomoGraphView (ours) 0.7527 0.6464 0.6711 0.2924
Kidney Tumors CT FMCIB 0.6828 0.6392 0.6785 0.2706
@@
Table 7: Quantitative dataset-specific metrics, showing median z-axis spacing, its corresponding z-axis spacing interquartile range (IQR) and z-axis alignment of the target anatomical structure. To quantify anatomical z-axis alignment, we apply principal component analysis (PCA) to the 3D segmentation masks to identify each structure’s major axis. The resulting alignment score is computed as the absolute cosine similarity between this major axis and the anatomical z-axis, yielding a value between 0 (no alignment) and 1 (perfect alignment). The unit measuring median slice spacing and IQR slice spacing is given in millimeters ($m$).
Metric Unit Breast Brain Head-Neck Kidney Liver Soft-Tissue
z-axis spacing (median) $m$ 1.00 1.00 2.00 3.00 1.00 4.80
z-axis spacing (IQR) $m$ 0.10 0.00 0.00 3.75 0.00 5.86
z-axis alignment n/a 0.40 0.32 0.92 0.51 0.10 0.93